# Complex potential energy surfaces: gradients with projected CAP technique

### v1

@@ 3. Computational details
SA-CASSCF geometry optimization of formaldehyde anion was performed with no symmetry constraints. Several state-averaging schemes have been used. In all cases, the state-averaging was performed over anionic states. The converged SA-CASSCF orbital were then used to perform CASCI calculation for the ground state of the neutral to get the reference energies. This choice of state-averaging scheme can introduce imbalance in description of the neutral and anion, which in turn can affect the computed resonance position. Yet, one can expect that not including the neutral in the state-averaging does not lead to detrimental description of the complex potential energy surface of the anionic resonance, and therefore, it’s equilibrium geometry. As the primary focus of this work is on nuclear gradients for the complex potential energy surfaces, this choice of methodology does not affect the main conclusions of the work. Two diffuse basis sets centered on the ghost atoms were used. The first basis set consisted of four -type basis functions (4p). The second basis included two -, five - and two - type even-tempered basis functions. In both cases, the initial exponents were chosen based on the most diffuse exponent in the parent basis. The scaling factor of two was used to generate the exponents of the even-tempered functions.<– not true always, we have used a combination of 1.5 and 2.0. The scaled exponents are available in the Supporting Information
@@ 4.2.2. Formaldehyde anion: shape resonance
The resonance parameters for the equilibrium geometry of the neutral and of the shape resonance are listed in the Table 3. Similarly to what has been observed for the N2, SA-CASSCF significantly overestimates resonance energy at the equilibrium geometry of the neutral placing the state at 2.4- 2.6 eV above the neutral. Including of the dynamic electron correlation effects brings the resonance energy down to 0.9 eV, which is in much better agreement with earlier more accurate theoretical estimates38, 79, 37, 80 and with the experimental reference value of 0.8681. Interestingly, the widths obtained with the three methods rather close (variation within 0.2 eV). Notably, while extension of the diffuse basis set from [4p] to [2s5p2d] leads to the shift in resonance energy of 0.2 eV and to the increase in the width of 0.08 eV, the effects on the optimized geometry are negligible (Table 2). Same trends for resonance energetic parameters, i.e. SA-CASSCF overestimating the resonance position and MR-CIS bringing the energy down, are observed for other molecular shape resonances, and, therefore, below we only focus on the equilibrium geometries rather than energetic parameters, which are reported in Supporting Information.

### v2

@@ 3. Computational details
SA-CASSCF geometry optimization of formaldehyde anion was performed with no symmetry constraints. Several state-averaging schemes have been used. In all cases, the state-averaging was performed over anionic states. The converged SA-CASSCF orbital were then used to perform CASCI calculation for the ground state of the neutral to get the reference energies. This choice of state-averaging scheme can introduce imbalance in description of the neutral and anion, which in turn can affect the computed resonance position. Yet, one can expect that not including the neutral in the state-averaging does not lead to detrimental description of the complex potential energy surface of the anionic resonance, and therefore, it’s equilibrium geometry. As the primary focus of this work is on nuclear gradients for the complex potential energy surfaces, this choice of methodology does not affect the main conclusions of the work. Two diffuse basis sets centered on the ghost atoms were used. The first basis set consisted of four -type basis functions (4p). The second basis included two -, five - and two - type even-tempered basis functions. In both cases, the initial exponents were chosen based on the most diffuse exponent in the parent basis. The scaled exponents are available in the Supporting Information
@@ 4.2.2. Formaldehyde anion: shape resonance
The resonance parameters for the equilibrium geometry of the neutral and of the shape resonance are listed in the Table S3. Similarly to what has been observed for the N2, SA-CASSCF significantly overestimates resonance energy at the equilibrium geometry of the neutral placing the state at 2.4- 2.6 eV above the neutral. Including of the dynamic electron correlation effects brings the resonance energy down to 0.9 eV, which is in much better agreement with earlier more accurate theoretical estimates38, 79, 37, 80 and with the experimental reference value of 0.8681. Interestingly, the widths obtained with the three methods rather close (variation within 0.2 eV). Notably, while extension of the diffuse basis set from [4p] to [2s5p2d] leads to the shift in resonance energy of 0.2 eV and to the increase in the width of 0.08 eV, the effects on the optimized geometry are negligible (Table 2). Same trends for resonance energetic parameters, i.e. SA-CASSCF overestimating the resonance position and MR-CIS bringing the energy down, are observed for other molecular shape resonances, and, therefore, below we only focus on the equilibrium geometries rather than energetic parameters, which are reported in Supporting Information.
@@ 6. Acknowledgments
Supporting Information
Complex potential energy surfaces: gradients with projected CAP technique
Soubhik Mondal and Ksenia B. Bravaya
Department of Chemistry, Boston University, Boston, Massachusetts, 02215, USA
S1. Comparison of CAP contributions to the gradient: CAP-EOM-EA-CCSD vs pCAP-EOM-EA-CCSD
| CAP-EOM-EA-CCSD | pCAP-EOM-EA-CCSD | |||||
| , : (2.605, 0.542) eV | , : (2.605, 0.546) eV | |||||
| Atom | X | Y | Z | X | Y | Z |
| N1 | 0.00000000 | 0.00000000 | 0.0567769 | 0.00000000 | -0.00000000 | 0.05810603 |
| N2 | 0.00000000 | 0.00000000 | -0.0567769 | 0.00000000 | -0.00000000 | -0.05810603 |
In Table S1, we report CAP contributions to the gradient for a single point calculation of resonance in CAP-EOM-EA-CCSD1 and pCAP-EOM-EA-CCSD2 levels respectively.
The similarities in magnitudes and signs of the CAP gradients from two different methods underlines that firstly, complex 1-RDMs from our method (S1) is similar to CAP-EOM-EA-CCSD where CAP is introduced in HF level and secondly, the evaluation of CAP derivative terms implementation on the grid is correct. We don’t expect the complex densities and CAP terms to be exactly same even though the complex energies are same because they are entirely different methods.
In ‘diag’ basis, we evaluate the 1-particle reduce density matrices (1-RDM) as following,
,where and ,where are the 1-TDM (transition density matrices) between state I and J in ‘diag’ and ZO basis respectively. If I and J are the same, they correspond to state density matrices.
S2. Rotation vector contribution to the gradient
We utilize the orthogonality relation of the rotation matrix : and the eigenvalue equation: , , where are the complex eigen-values. It follows that,
S3. Tracking resonance state: using attachment and detachment densities
The tracking algorithm we deploy when full wave-function overlap is not available, is through comparison of attachment and detachment densities3, 4 in subsequent steps, an approach which is also adopted in the Q-CHEM program5. Equation-wise, the overlap metric between state I and J from consecutive steps is evaluated as following4,
where, and are attachment and detachment densities in ‘diag’ basis for old and new iterations respectively. In equation (S3), the norm of the matrix M is evaluated as with being eigenvalues of . Essentially, equation (S3) provides an estimate of how the density of a state changes between two iterations, if the densities are close, is 1.
The attachment and detachment densities for state are calculated using prescriptions from reference 3 and the one particle difference density matrix ()3 is built with the reference state chosen as the neutral state evaluated in the same orthogonal basis to that of the correlated anionic states. The density contraction to ‘diag’ from ZO basis is done using equation (S1).
S4. pCAP parameters for initial calculations at equilibrium geometries of the neutral & energetics
| Methods/basis | ||||
| SA-10-CASSCF(5e, 13o) | 2.77 | 2.77 | 4.88 | 0.00790 |
| cc-pVTZ+[2s5p2d] | ||||
| MR-CIS/SA-6-CASSCF(7e, 10o) | 2.77 | 2.77 | 4.88 | 0.00854 |
| cc-pVTZ+[2s5p2d] | ||||
| MR-CIS/SA-6-CASSCF(5e, 9o) | 2.77 | .2.77 | 4.88 | 0.00724 |
| cc-pVTZ+[2s5p2d] | ||||
| MR-CIS/SA-10-CASSCF(5e, 8o) | 2.76 | 2.76 | 4.88 | 0.00826 |
| aug-cc-pVTZ+X[3s3p3d] |
| Methods/basis | ||||||
| SA-15-CASSCF(3e, 16o)/ | 5.98 | 2.426 | 0.338 | 0.01656 | 1.281 | 0.232 |
| cc-pVTZ+X[4p] | 3.88 | |||||
| 2.93 | ||||||
| SA-19-CASSCF(3e, 20o)/ | 5.99 | 2.633 | 0.416 | 0.02888 | 1.086 | 0.687 |
| cc-pVTZ+X[2s5p2d] | 3.89 | |||||
| 2.94 | ||||||
| MRCIS/SA-7-CASSCF(3e, 8o)/ | 6.05 | 0.931 | 0.211 | 0.00196 | 0.168 | 0.101 |
| cc-pVDZ+[3p] (Cs symmetry) | 3.94 | |||||
| 2.97 | ||||||
| Methods/basis | ||||||
| SA-15-CASSCF(3e, 20o)/ | 7.62 | 3.119 | 0.634 | 0.01772 | 2.350 | 0.364 |
| cc-pVTZ+X[3p] | 4.20 | |||||
| 6.02 | ||||||
| MR-CIS/SA-6-CASSCF(3e, 7o)/ | 7.14 | 2.617 | 0.592 | 0.00465 | 2.079 | 0.348 |
| cc-pVDZ+[3p] ( symmetry) | 3.50 | |||||
| 4.69 | ||||||
| Methods/basis | ||||||
| SA-16-CASSCF(3e, 17o)/ | 9.45 | 3.889 | 0.199 | 0.00743 | 1.591 | 0.258 |
| cc-pVTZ+X[2s5p2d] | 5.29 | |||||
| 3.56 | ||||||
| SA-16-CASSCF(3e, 17o)/ | 9.45 | 4.116 | 0.240 | 0.00864 | 1.580 | 1.125 |
| cc-pVTZ+X[3p] | 5.28 | |||||
| 3.52 | ||||||
| SA-15-CASSCF(3e, 23o)/ | 9.54 | 3.739 | 0.065 | 0.04005 | 1.636 | 0.103 |
| cc-pVTZ+X[4p] | 6.36 | |||||
| 6.33 | ||||||
| SA-15-CASSCF(3e, 23o)/ | 9.54 | 4.160 | 0.133 | 0.03878 | 1.674 | 0.496 |
| cc-pVTZ+X[5p] | 5.29 | |||||
| 3.56 | ||||||
S5. Augmented diffuse basis
The main parent basis used in our calculations is cc-pVTZ, unless mentioned this is the case for the following.
S5.1.
2s,2d scaling exponent of . 5p scaling exponent of .
ΨNΨΨΨ0 ΨS 1 1.00 Ψ0.0893500 1.0000000 ΨS 1 1.00 Ψ0.0446750 1.0000000 ΨP 1 1.00 Ψ0.1150 1.0000000 ΨP 1 1.00 Ψ0.0766667 1.0000000 ΨP 1 1.00 Ψ0.0511111 1.0000000 ΨP 1 1.00 Ψ0.0340741 1.0000000 ΨP 1 1.00 Ψ0.0227160 1.0000000 ΨD 1 1.00 Ψ0.234500 1.000000 ΨD 1 1.00 Ψ0.117250 1.000000
3s,3p.3d scaling exponent for for aug-cc-pVTZ.
ΨXΨΨ0 ΨS 1 1.00 Ψ0.0288 1.00000000 ΨS 1 1.00 Ψ0.0144 1.00000000 ΨS 1 1.00 Ψ0.0072 1.00000000 ΨP 1 1.00 Ψ0.02455 1.0000000Ψ ΨP 1 1.00 Ψ0.012275 1.00000000 ΨP 1 1.00 Ψ0.0061375 1.00000000 ΨD 1 1.00 Ψ0.0755 1.000000 ΨD 1 1.00 Ψ0.03775 1.00000 ΨD 1 1.00 Ψ0.018875 1.00000
S5.2.
3p scaling exponent with cc-pVDZ.
ΨCΨΨ0 ΨP 1 1.00 Ψ0.13765 1.0 ΨP 1 1.00 Ψ0.068825 1.0 ΨP 1 1.00 Ψ0.0344125 1.0
4p scaling exponent
ΨX 0 ΨP 1 1.00 Ψ0.0806 1.000000D+00 ΨP 1 1.00 Ψ0.05373 1.000000D+00 ΨP 1 1.00 Ψ0.03582 1.000000D+00 ΨP 1 1.00 Ψ0.02388Ψ1.00
2s,2d scaling exponent of . 5p scaling exponent of .
ΨX 0 ΨS 1 1.00 Ψ0.05135Ψ1.0 ΨS 1 1.00 Ψ0.025675 1.0000000 ΨP 1 1.00 Ψ0.0806 1.000000D+00 ΨP 1 1.00 Ψ0.05373 1.000000D+00 ΨP 1 1.00 Ψ0.03582 1.000000D+00 ΨP 1 1.00 Ψ0.02388Ψ1.00 ΨP 1 1.00 Ψ0.01592Ψ1.00 ΨD 1 1.00 Ψ0.159 1.0000000 ΨD 1 1.00 Ψ0.0795 1.0000000
S5.3.
3p scaling exponent with cc-pVDZ.
ΨP 1 1.00 Ψ0.13765 1.0 ΨP 1 1.00 Ψ0.068825 1.0 ΨP 1 1.00 Ψ0.0344125 1.0
3p scaling exponent .
ΨCΨ0 ΨP 1 1.00 Ψ0.0806Ψ1. ΨP 1 1.00 Ψ0.05373333333333333Ψ1. ΨP 1 1.00 Ψ0.03582222222222222Ψ1.
S5.4.
4p/5p: scaling exponent of .
X 0 P 1 1.00 0.06045 1.000000D+00 P 1 1.00 0.030225 1.000000D+00 P 1 1.00 0.0151125 1.000000D+00 P 1 1.00 0.00755625Ψ 1.000000D+00 P 1 1.00 0.003778125 1.000000D+00
3p, scaling exponent .
X 0 P 1 1.00 0.0806 1.000000D+00 P 1 1.00 0.05373 1.000000D+00 P 1 1.00 0.03582 1.000000D+00
2s,2d scaling exponent is ). 5p scaling exponent is ..
X 0 S 1 1.00 0.05135Ψ1.0 S 1 1.00 0.025675 1.0000000 P 1 1.00 0.0806 1.000000D+00 P 1 1.00 0.05373 1.000000D+00 P 1 1.00 0.03582 1.000000D+00 P 1 1.00 0.02388Ψ1.0 P 1 1.00 0.01592Ψ1.0 D 1 1.00 0.159 1.0000000 D 1 1.00 0.0795 1.0000000
S6. Initial geometries
S6.1. Initial geometry (in Å) of with cc-pVTZ basis
with symmetry.
N 0.00000000 0.00000000 0.54743386 N 0.00000000 0.00000000 -0.54743386
S6.2. Initial geometry (in Å) of with cc-pVTZ basis
with symmetry.
C 0.539081 -0.009227 0.000000 O -0.636180 -0.086613 0.000000 H 1.181355 -0.893267 0.000000 H 1.056513 0.968385 0.000000
S6.3. Initial geometry (in Å) of with cc-pVDZ basis
with symmetry.
C 0.535543212514 -0.008891235494 0.00000000 O -0.643690629882 -0.088924526013 0.00000000 H 1.185776060252 -0.897982505052 0.00000000 H 1.063138712502 0.975076216165 0.00000000
S6.4. Initial geometry (in Å) of with cc-pVTZ basis
with symmetry.
C 0.150321 -0.378974 0.000000 H 0.213778 -1.461673 0.000000 O 1.082107 0.341490 0.000000 O -1.109733 0.010919 0.000000 H -1.143257 0.957261 0.000000
S6.5. Initial geometry (in Å) of with cc-pVDZ basis
with symmetry.
C 0.151159 -0.375236 0.000000 H 0.213978 -1.480920 0.000000 O 1.079429 0.352747 0.000000 O -1.113650 0.010337 0.000000 H -1.137699 0.962095 0.000000
S6.6. Initial geometry (in Å) of with cc-pVTZ basis
with symmetry.
C -0.0179750928 -0.1136932028 -0.0115175428 H -0.0496468803 0.0145973544 1.0544147734 H 0.9369204904 0.0146037045 -0.4862662659 C -1.1084247201 -0.4170070183 -0.7097827982 H -1.0767529326 -0.5452975755 -1.7757145852 H -2.0633203033 -0.5453028673 -0.2350335460
S6.7. Initial distorted geometry (in Å) of with cc-pVTZ basis
with symmetry.
ΨC 0.670519 -0.000015 0.000000 ΨC -0.670519 0.000015 0.000000 ΨH 1.238290 0.000271 0.922918 ΨH -1.238290 -0.000271 0.922918 ΨH 1.238290 0.000271 -0.922918 ΨH -1.238290 -0.000271 -0.922918
S6.8. Initial distorted geometry (in Å) of with cc-pVDZ basis
with symmetry.
ΨC 0.670518 0.000007 0.000000 ΨC -0.670518 -0.000007 0.000000 ΨH 1.238283 -0.000164 0.922917 ΨH -1.238283 0.000164 0.922917 ΨH 1.238283 -0.000164 -0.922917 ΨH -1.238283 0.000164 -0.922917